In a large scavenger hunt, the number of people on each team is distributed as follows:...

  1. In a large scavenger hunt, the number of people on each team is distributed as follows:

    A team is selected.

    1. a) What is the probability that there is only one person on the team?

    2. b) What is the probability there are between 3 and 6 people on the team?

    3. c) What is the average number of people on a team?

    4. d) Determine the standard deviation.

x. 2. 4. 5. 6. 8

P(X=x). 0.15. 0.44. 0.26. 0.10 0.05
Answer #1

a) probability that there is only one person on the team =P(X=1)=0

(as from given distribution x=1 does not exist)

b) probability there are between 3 and 6 people on the team =P(3<=X<=6)=P(X=5)+P(X=6)=0.26+0.1

=0.36

c)

x P(x) xP(x) x2P(x)
2 0.15 0.300 0.600
4 0.44 1.760 7.040
5 0.26 1.300 6.500
6 0.1 0.600 3.600
8 0.05 0.400 3.200
total 4.360 20.940
E(x) =μ= ΣxP(x) = 4.3600
E(x2) = Σx2P(x) = 20.9400
Var(x)=σ2 = E(x2)-(E(x))2= 1.930
std deviation=         σ= √σ2 = 1.389

average number of people on a team =4.36

d)

standard deviation =1.389